OpenStudy (anonymous):

is the line passing through points a(-5,20 and b(2,-2( perpendicular to the line passing through points C(0,5) and D(4,12) a. yes their slopes are different b. no their slopes are not opposite reciprocals c. no their slopes are not the same d. yes their slopes are opposite reciprocals

OpenStudy (cwrw238):

calculate the slopes of the 2 lines if they are perpendicular then m1* m2 = -1 (m1 and m2 being the 2 slopes)

OpenStudy (rajat97):

@cwrw238 is right find the slopes by the formula y2-y1/x2-x1 so you may get the answers a,c and d

OpenStudy (anonymous):

i still dont understand it

OpenStudy (rajat97):

let the point (2,-2)be (x1,y1) and let the point (-5,2)be (x2,y2) thus by putting the values of x1,y1,x2 and y2 in the specified formula, you may get the slope of the first line do the same with the other two points let (0,5) be (x1,y1) and let (4,12) be (x2,y2) and put the values in the formula for slope(specified above) then multiply the slopes of both the lines. If the product comes out to be -1, the lines are perpendicular to each other!

OpenStudy (anonymous):

u dont understand cuz u stupid.

OpenStudy (rajat97):

rude girl @laylaloveeee

OpenStudy (cwrw238):

very rude

OpenStudy (anonymous):

You're going to want to find the slope of C and D and A and B, because a line perpendicular to another line will have the negative reciprocal slope.... Slope fomula is Y2-Y1 divided by X2-X1

OpenStudy (cwrw238):

heres the slope of the first line done for you y1 = 20, y2 = -2 x1 = -5, x2 = 2 so slope = -2-20 / 2- (-5) = -22 / 10 now try working out the slope of 2nd line in the same way